# On the Identification of Temporally Causal Representation with Instantaneous Dependence

### v2

@@ 5.1.2 Results and Discussion
Quantitative Results: Experiment results of the simulation datasets are shown in Table 1. According to the experiment results, we can obtain the following conclusions: 1) we can find that the proposed IDOL model achieves the highest MCC performance, reflecting that our method can identify the latent variables under the temporally latent process with instantaneous dependency. 2) our IDOL method also outperforms G-CaRL and iCITRIS, which leverage grouping and interventions for identifiability of latent variables with instantaneous dependencies, reflecting that our method does not require a stricter assumption for identifiability. 3) Compared with the existing methods for temporal data like TDRL, PCL, and TCL, the experiment results show that our method can identify the general stationary latent process with instantaneous dependencies. 4) Considering the results of our IDOL method of Dataset A, B, and C with different levels of sparsity, we can find that the more sparse the latent process is, the better identifiable performance the model can achieve, which indirectly reflects the necessity of the sparsity assumption. 5) According to the results of the Dense dataset, the IDOL cannot well identify the latent variables, validating the boundary of our theoretical results.
Table 1: Experiments results on simulation data.
@@
Qualitative Results: Figure 4 provides visualization results on dataset B of the latent causal process of IDOL, TDRL, G-CaRL, and iCITRIS. Note that the visualization results of our method are generated from $\mathbf{J}_{d}$ and $\mathbf{J}_{e}$. According to the experiment results, we can find that the proposed method can well reconstruct the latent causal relationship, which validates the theoretical results. Moreover, since TDRL employs the conditional independent assumption, it does not reconstruct any instantaneous causal relationships. Besides, without any accurate grouping or intervention information, the G-CaRL and iCITRIS can not reconstruct the correct latent causal relationships. Please refer to Appendix A4.1.4 for more details.
@@ 5.2.1 Experiment Setup
Datasets. To evaluate the effectiveness of our IDOL method in real-world scenarios, we conduct experiments on two human motion datasets: Human3.6M [43] and HumanEva-I [44], which record the location and orientation of local coordinate systems at each joint. We consider these datasets since the joints can be considered as latent variables and they naturally contain instantaneous dependencies. We choose several motions to conduct time series forecasting. For the Human3.6M datasets, we consider 4 motions: Gestures (Ge), Jog (J), CatchThrow (CT), and Walking (W). For HumanEva-I dataset, we consider 6 motions: Discussion (D), Greeting (Gr), Purchases (P), SittingDown (SD), Walking (W), and WalkTogether (WT). For each dataset, we select several motions and partition them into training, validation, and test sets. Please refer the Appendix A4.3 for the dataset descriptions.
Baselines. To evaluate the effectiveness of the proposed IDOL, we consider the following state-of-the-art deep forecasting models for time series forecasting. First, we consider the conventional methods for time series forecasting including Autoformer [45], TimesNet [4] and MICN [46]. Moreover, we consider several latest methods for time series analysis like CARD [47], FITS [48], and iTransformer [49]. Finally, we consider the TDRL [39]. We repeat each experiment over 3 random seeds and publish the average performance. Please refer to Appendix A4.2 for more experiment details.
@@ 5.2.2 Results and Discussion
Figure 5: The illustration of visualization of different methods of Walking. The red lines denote the ground-truth motions, those lines in other colors denote the prediction motions of different methods.
Quantitative Results: Experiment results of the real-world datasets are shown in Table LABEL:tab:main_exp. More experiment results are shown in Appendix A4.4. According to the experiment results, our IDOL model significantly outperforms all other baselines on most of the human motion forecasting tasks. Specifically, our method outperforms the most competitive baseline by a clear margin of 4%-34% and promotes the forecasting accuracy substantially on complex motions like the SittingDown (SD) and the Walking (W). This is because the human motion datasets contain more complex patterns described by stable causal structures, and our method can learn the latent causal process with identifiability guarantees. It is noted that our method achieves a better performance than that of TDRL, which does not consider the instantaneous dependencies among latent variables. We also find that our model achieves a comparable performance in the motion of CatchThrow (CT), this is because the size of this dataset is too small for the model to model the distribution of the observations. We further consider a more complex synthetic mixture of human motion forecasting, experiment results are shown in Appendix A4.4. Please refer to Appendix A4.5 for the experiment results of ablation studies.
Qualitative Results: To further show the effectiveness of the proposed method, we also visualize the forecasting results of different models. Visualization results in Figure 5 illustrate how the forecasting results align the ground-truth motions, where the red lines denote the ground-truth motions and the lines in other colors denote the forecasting results of different methods. According to the experiment results, we can find that the forecasting results of our method achieve the best alignment of the ground truth. In the meanwhile, the other methods that do not consider the instantaneous dependencies of latent variables, may generate some exaggerated body movements, for example, the hyperflexed knee joint and the actions that violate the physics rules, which reflects the importance of considering the instantaneous dependencies of latent causal process of time series data.
@@ 6 Conclusion
This paper proposes a general identifiability framework for temporal causal representation learning with instantaneous dependencies. Different from existing methods that require the assumptions of grouping observations and interventions, the proposed IDOL model employs the sparse latent process assumption, which is easy to satisfy in real-world time series data. We also devise a variational-inference-based method with sparsity regularization to build the gap between theories and practice. Experiment results on simulation datasets evaluate the effectiveness of latent variables identification and latent directed acyclic graphs reconstruction. Evaluation in human motion datasets with instantaneous dependencies reflects the practicability in real-world scenarios.
@@
[…]
1
Appendix A1 Related Works
A1.1 Identifiability of Causal Representation Learning
@@
To achieve the causal representation [50, 51, 52] for time series data, several researchers leverage the independent component analysis (ICA) to recover the latent variables with identification guarantees [53, 54, 55, 56]. Conventional methods assume a linear mixing function from the latent variables to the observed variables [20, 21, 57, 22]. To relax the linear assumption, researchers achieve the identifiability of latent variables via nonlinear ICA by using different types of assumptions like auxiliary variables or sparse generation process [27, 58, 59, 60, 61]. As for the methods that employ the auxiliary variables, Aapo et al. first achieve the identifiability by assuming the latent sources with exponential family and introducing auxiliary variables e.g., domain indexes, time indexes, and class labels[24, 18, 30, 31]. To further relax the exponential family assumption, Zhang et al. [62, 25, 63, 64, 25] achieve the component-wise identification results for nonlinear ICA with a $12n+$ 1 number of auxiliary variables for $n$ latent variables. Recently, Li et al. [65] further relax to $n+1$ number of auxiliary variables and achieve subspace identifiability. To seek identifiability in an unsupervised manner, researchers employ the assumption of structural sparsity to achieve identifiability [32, 26, 27, 66]. Specifically, Lachapelle et al. [67, 68] proposed mechanism sparsity regularization as an inductive bias to identify the causal latent factors and achieve the identifiability on several scenarios like multi-task learning. They further show the identifiablity results up to the consistency relationship [69], which allows the partial disentanglement of latent variables. Recently, Zhang et al. [36] use the sparse structures of latent variables to achieve identifiability under distribution shift. Different from these methods that assume sparsity exists in the generation process between the latent source and the observation, we assume the latent process sparsity, where sparsity exists in the transition of latent variables and is common in real-world time series data.dependencies.
A1.2 Nonlinear ICA for Time Series Data
@@
Our work is also related to the temporally causal representation. Existing methods for temporally causal representation [64, 70, 28, 71] usually rely on the conditional independence assumption, where the latent variables are mutually independent conditional on their time-delayed parents. Specifically, Aapo et al. [18] leverage the independent sources principle and the variability of data segments to achieve identifiability on nonstationary time series data. They further use permutation-based contrastive [30] to achieve identifiability on stationary time series data. The recent advancements in temporally causal representation include LEAP [23], TDRL [39], which leverage the properties of independent noises and variability historical information. However, this assumption is too strong to meet in real-world scenarios since instantaneous dependencies are common in real-world scenarios like human motion. To solve this problem, researchers introduce further assumptions. For example, Lippe et al. [34] propose the i-Citris, which demonstrates identifiability by assuming interventions on latent variables. And Morioka et al. [35] propose G-CaRL, which yields identifiability with the assumption of the grouping of observational variables. However, these assumptions are still hard to meet in practice. Hence we propose a more relaxed assumption, transition sparsity, to achieve identifiability under time series data with instantaneous
A1.3 Causal Discovery with Latent Variables
@@
Several studies are proposed to discover causally related latent variables [72, 73, 74, 75, 76]. Specifically, Huang et al. [77, 78] leverage the vanishing Tetrad conditions or rank constraints to identify latent variables in linear-Gaussian models. Cai et al. [79, 80] further draw upon non-Gaussianity in their analysis for linear, non-Gaussian scenarios. Other methods aim to reconstruct the hierarchical structures of the latent variables, like [79, 77]. However, these methods usually use the linear assumption and can hardly handle the real-world time series data with complex nonlinear relationships.
A1.4 Instantaneous Dependencey of Time Series Data
@@
When the discrete time series data is sampled in a low-frequency or low-resolution manner, instantaneous dependence [81, 82, 83, 84, 85] occurs where variables at each time step are not independent given their historical observations. [86] first discuss the instantaneous dependence from the perspective of noise. Lots of works investigate instantaneous dependency from the causal view. Specifically, Hyvärinen et al. [87] estimate both instantaneous and lagged effects via a non-Gaussian model. Gong et al. [88, 89] discover causal structure with instantaneous dependency from subsampled and aggregated data. Recently, Zhu et al. [90] consider instantaneous dependency in reinforcement learning.
A1.5 Time Series Forecasting
@@
Time series forecasting [91, 92] is one of the most popular research problems in recent years. Recently, the deep-learning-based methods have made great progress, which can be categorized according to different types of neural architectures like the RNN-based methods [93, 94, 95], CNN-based models [96, 46, 4], and the methods based on state-space model [97, 98, 99]. Recently, Transformer-based methods [100, 45, 101] further push the development of time series forecasting. However, these methods seldom consider the instantaneous dependencies of time series data.
Appendix A2 Proof of Theory
A2.1 Relationships between Ground-truth and Estimated Latent Variables
@@
For a series of observations $\mathbf{x}_{t}\in\mathbb{R}^{n}$ and estimated latent variables $\mathbf{\hat{z}}_{t}\in\mathbb{R}^{n}$ with the corresponding process $\hat{f}_{i},\hat{p}(\epsilon),\hat{g}$, where $\hat{g}$ is invertible, suppose the process subject to observational equivalence $\mathbf{x}_{t}=\mathbf{\hat{g}}(\mathbf{\hat{z}}_{t})$. Let $\mathbf{c}_{t}\triangleq\{\mathbf{z}_{t-1},\mathbf{z}_{t}\}$ and $\mathcal{M}_{\mathbf{c}_{t}}$ be the variable set of 2 consecutive timestamps and the corresponding Markov network respectively. Suppose assumptions
  • A1 (Smooth and Positive Density): The probability function of the latent variables ${\mathbf{c}}_{t}$ is smooth and positive, i.e., $p_{{\mathbf{c}}_{t}}$ is third order differentiable and $p_{{\mathbf{c}}_{t}}>0$ over $\mathbb{R}^{2n}$,
@@
    $\small\begin{split}w(m)=&\Big{(}\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,1}^{2}\partial z_{t-2,m}},\cdots,\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,2n}^{2}\partial z_{t-2,m}}\Big{)}\oplus\\ &\Big{(}\frac{\partial^{2}\log p(c_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,1}\partial z_{t-2,m}},\cdots,\frac{\partial^{2}\log p(c_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,2n}\partial z_{t-2,m}}\Big{)}\oplus\Big{(}\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}\partial c_{t,j}\partial z_{t-2,m}}\Big{)}_{(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{c}}_{t}})},\end{split}$ (13)
    where $\oplus$ denotes concatenation operation and $(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{c}}_{t}})$ denotes all pairwise indice such that $c_{t,i},c_{t,j}$ are adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$. For $m\in[1,\cdots,n]$, there exist $4n+2|\mathcal{M}_{{\mathbf{c}}_{t}}|$ different values of ${\mathbf{z}}_{t-2,m}$, as the $4n+2|\mathcal{M}_{{\mathbf{c}}_{t}}|$ values of vector functions $w(m)$ are linearly independent,
hold true. Then for any two different entries $\hat{c}_{t,k},\hat{c}_{t,l}\in\hat{{\mathbf{c}}}_{t}$ that are not adjacent in the Markov network $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$ over estimated $\hat{{\mathbf{c}}}_{t}$,
(i) Each ground-truth latent variable $c_{t,i}\in{\mathbf{c}}_{t}$ is a function of at most one of $\hat{c}_{t,k}$ and $\hat{c}_{t,l}$,
(ii) For each pair of ground-truth latent variables $c_{t,i}$ and $c_{t,j}$ that are adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$ over ${\mathbf{c}}_{t}$, they can not be a function of $\hat{c}_{t,k}$ and $\hat{c}_{t,l}$ respectively.
@@
$\small p(\hat{{\mathbf{x}}}_{t})=p({\mathbf{x}}_{t})\Longleftrightarrow p(\hat{g}(\hat{{\mathbf{z}}}_{t}))=p(g(z_{t}))\Longleftrightarrow p((g^{-1}\circ\hat{g})(\hat{{\mathbf{z}}}_{t}))=p({\mathbf{z}}_{t})\Longleftrightarrow p(h_{z}(\hat{{\mathbf{z}}}_{t}))=p({\mathbf{z}}_{t}),$ (14)
where $\hat{g}:\mathcal{Z}\rightarrow\mathcal{X}$ denotes the estimated mixing function, and $h:=g^{-1}\circ\hat{g}$ is the transformation between the ground-truth latent variables and the estimated ones. Since $\hat{g}$ and $g$ are invertible, $h$ is invertible as well. Since Equation 14 holds true for all time steps, there must exist an invertible function $h_{c}$ such that $p(h_{c}(\hat{{\mathbf{c}}}_{t}))=p({\mathbf{c}}_{t})$, whose Jacobian matrix at time step $t$ is
$\mathbf{J}_{h_{c},t}=\begin{bmatrix}\mathbf{J}_{h_{z},t-1}&0\\ 0&\mathbf{J}_{h_{z},t}\end{bmatrix}.$ (15)
@@
where * denotes any matrix. Since ${\mathbf{x}}_{t-2}$ can be fully characterized by itself, the left top and right top block are 1 and 0 respectively, and the determinant of this Jacobian matrix is the same as $|\mathbf{J}_{h_{c},t}|$. Therefore, we have:
$p(\hat{{\mathbf{c}}}_{t},{\mathbf{x}}_{t-2})=p({\mathbf{c}}_{t},{\mathbf{x}}_{t-2})|\mathbf{J}_{h_{c},t}|.$ (16)
Dividing both sides of Equation (7) by $p({\mathbf{x}}_{t-2})$, we further have:
$p(\hat{{\mathbf{c}}}_{t}|{\mathbf{x}}_{t-2})=p({\mathbf{c}}_{t}|{\mathbf{x}}_{t-2})|\mathbf{J}_{h_{c},t}|.$ (17)
Since $p({\mathbf{c}}_{t}|{\mathbf{x}}_{t-2})=p({\mathbf{c}}_{t}|g({\mathbf{z}}_{t-2}))=p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})$, and similarly $p(\hat{{\mathbf{c}}}_{t}|{\mathbf{x}}_{t-2})=p(\hat{{\mathbf{c}}}_{t}|\hat{\mathbf{z}}_{t-2})$, we have:
$\log p(\hat{{\mathbf{c}}}_{t}|\hat{\mathbf{z}}_{t-2})=\log p({{\mathbf{c}}}_{t}|{\mathbf{z}}_{t-2})+\log|\mathbf{J}_{h_{c},t}|.$ (18)

### v3

@@ 5.1.2 Results and Discussion
Quantitative Results: Experiment results of the simulation datasets are shown in Table 2. The proposed IDOL model achieves the highest MCC performance, reflecting that our method can identify the latent variables under the temporally latent process with instantaneous dependency. According to the experiment results, we can obtain the following conclusions: 1) our IDOL method also outperforms G-CaRL and iCITRIS, which leverage grouping and interventions for identifiability of latent variables with instantaneous dependencies, reflecting that our method does not require a stricter assumption for identifiability. 2) Compared with the existing methods for temporal data like TDRL, PCL, and TCL, the experiment results show that our method can identify the general stationary latent process with instantaneous dependencies. 3) According to the results of the dense dataset E, the IDOL cannot well identify the latent variables, validating the boundary of our theoretical results. 4) We further analyze the effectiveness of the sparsity regularization. Please refer to Appendix F.2.4,F.1.4 for more details.
Table 2: Experiments results of MCC on simulation data. The iCITRIS algorithm encountered an out-of-memory (OOM) issue when running on dataset F due to the high dimensionality of the latent variables, which caused the code to fail to execute. This is indicated by the symbol ’-’.
@@
Qualitative Results: Figure 4 provides visualization results on dataset A of the latent causal process of IDOL, TDRL, G-CaRL, and iCITRIS. Note that the visualization results of our method are generated from $\mathbf{J}_{d}$ and $\mathbf{J}_{e}$. According to the experiment results, we can find that the proposed method can reconstruct the latent causal relationship well, which validates the theoretical results. Please note that here, not only the Markov equivalence class but also the causal graph can be identified for dataset A, as shown in Figure 4. Please refer to Appendix B.5 for more details. Moreover, since TDRL employs the conditional independent assumption, it does not reconstruct any instantaneous causal relationships. Besides, without any accurate grouping or intervention information, the G-CaRL and iCITRIS can not reconstruct the correct latent causal relationships. Please refer to Appendix E.3.1 for more details.
@@ 5.2.1 Experiment Setup
Datasets. To evaluate the effectiveness of our IDOL method in real-world scenarios, we conduct experiments on two human motion datasets: Human3.6M (6682899) and HumanEva-I (DBLP:journals/ijcv/SigalBB10), which record the location and orientation of local coordinate systems at each joint. We consider these datasets since the joints can be considered as latent variables and they naturally contain instantaneous dependencies. We choose several motions to conduct time series forecasting. For the Human3.6M datasets, we consider 4 motions: Gestures (Ge), Jog (J), CatchThrow (CT), and Walking (W). For HumanEva-I dataset, we consider 6 motions: Discussion (D), Greeting (Gr), Purchases (P), SittingDown (SD), Walking (W), and WalkTogether (WT). For each dataset, we select several motions and partition them into training, validation, and test sets. Please refer the Appendix F.2.1 for the dataset descriptions.
Baselines. To evaluate the effectiveness of the proposed IDOL, we consider the following state-of-the-art deep forecasting models for time series forecasting. First, we consider the conventional methods for time series forecasting including Autoformer (wu2021autoformer), TimesNet (wu2022timesnet) and MICN (wang2022micn). Moreover, we consider several latest methods for time series analysis like CARD (wang2023card), FITS (xu2024fits), and iTransformer (liu2024itransformer). Finally, we consider the TDRL (yao2022temporally). We repeat each experiment over 3 random seeds and publish the average performance. Please refer to Appendix F.2 for more experiment details.
@@ 5.2.2 Results and Discussion
Quantitative Results: Experiment results of the real-world datasets are shown in Table 3. We also conduct the Wilcoxon signed-rank test richard2021concepts on the reported accuracies, our method significantly outperforms the baselines, with a p-value threshold of 0.05. According to the experiment results, our IDOL model significantly outperforms all other baselines on most of the human motion forecasting tasks. Specifically, our method outperforms the most competitive baseline by a clear margin of 4%-34% and promotes the forecasting accuracy substantially on complex motions like the SittingDown (SD) and the Walking (W). This is because the human motion datasets contain more complex patterns described by stable causal structures, and our method can learn the latent causal process with identifiability guarantees. It is noted that our method achieves a better performance than that of TDRL, which does not consider the instantaneous dependencies among latent variables. We also find that our model achieves a comparable performance in the motion of CatchThrow (CT), this is because the size of this dataset is too small for the model to model the distribution of the observations. We further consider a more complex synthetic mixture of human motion forecasting, experiment results are shown in Appendix F.2.2. Please refer to Appendix F.2.4 for the experiment results of ablation studies. Moreover, we also investigate the proposed IDOL model on other high-dimension datasets in Appendix H.1.
@@ 6 Conclusion
This paper proposes a general framework for time series data with instantaneous dependencies to identify the latent variables and latent causal relations up to the Markov equivalence class. Furthermore, with mild assumption, the causal graph is also identifiable. Different from existing methods that require the assumptions of grouping observations and interventions, the proposed IDOL model employs the sparse latent process assumption, which is easy to satisfy in real-world time series data. We also devise a variational-inference-based method with sparsity regularization to build the gap between theories and practice. Experiment results on simulation datasets evaluate the effectiveness of latent variables identification and latent directed acyclic graphs reconstruction. Evaluation in human motion datasets with instantaneous dependencies reflects the practicability in real-world scenarios. There are two main limitations in our work. First, our method is not for high-dimensional time series data because modeling the complex dependencies through sparsity constraints is challenging. Second, our method relies on the invertible mixing process, but it may not hold in real-world scenarios. How to address these limitations and make our method more scalable will be an interesting direction.
7 Acknowledgment
We would like to acknowledge the support from NSF Award No. 2229881, AI Institute for Societal Decision Making (AI-SDM), the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Quris AI, Florin Court Capital, and MBZUAI-WIS Joint Program. Moreover, this research was supported in part by the National Science and Technology Major Project (2021ZD0111501), Natural Science Foundation of China (U24A20233).
@@ 7 Acknowledgment
Appendix A Related Works
A.1 Identifiability of Causal Representation Learning
@@
To achieve the causal representation (rajendran2024learning; mansouri2023object; wendong2024causal) for time series data, several researchers leverage the independent component analysis (ICA) to recover the latent variables with identification guarantees (yao2023multi; scholkopf2021toward; Liu2023CausalTriplet; gresele2020incomplete). Conventional methods assume a linear mixing function from the latent variables to the observed variables (comon1994independent; hyvarinen2013independent; lee1998independent; zhang2007kernel). To relax the linear assumption, researchers achieve the identifiability of latent variables via nonlinear ICA by using different types of assumptions like auxiliary variables or sparse generation process (zheng2022identifiability; hyvarinen1999nonlinear; hyvarinen2023identifiability; khemakhem2020ice; li2023identifying). As for the methods that employ the auxiliary variables, khemakhem2020variational first achieve the identifiability by assuming the latent sources with exponential family and introducing auxiliary variables e.g., domain indexes, time indexes, and class labels (khemakhem2020variational; hyvarinen2016unsupervised; hyvarinen2017nonlinear; hyvarinen2019nonlinear). To further relax the exponential family assumption, kong2022partial; xie2022multi; kong2023identification; yan2023counterfactual; xie2022multi achieve the component-wise identification results for nonlinear ICA with a $2n+1$ number of auxiliary variables for $n$ latent variables. Recently, li2024subspace further relax to $n+1$ number of auxiliary variables and achieve subspace identifiability. To seek identifiability in an unsupervised manner, researchers employ the assumption of structural sparsity to achieve identifiability (ng2024identifiability; lachapelle2022disentanglement; zheng2022identifiability; xu2024sparsity). Specifically, lachapelle2023synergies; lachapelle2022partial proposed mechanism sparsity regularization as an inductive bias to identify the causal latent factors and achieve the identifiability on several scenarios like multi-task learning. They further show the identifiability results up to the consistency relationship (lachapelle2024nonparametric), which allows the partial disentanglement of latent variables. Recently, (zhang2024causal) use the sparse structures of latent variables to achieve identifiability under distribution shift. Different from these methods that assume sparsity exists in the generation process between the latent source and the observation, we assume the sparse latent process, where sparsity exists in the transition of latent variables and is common in real-world time series data.
A.2 Nonlinear ICA for Time Series Data
@@
Our work is also related to the temporally causal representation. Existing methods for temporally causal representation (yan2023counterfactual; huang2023latent; halva2020hidden; lippe2022citris) usually rely on the conditional independence assumption, where the latent variables are mutually independent conditional on their time-delayed parents. Specifically, hyvarinen2016unsupervised leverage the independent sources principle and the variability of data segments to achieve identifiability on nonstationary time series data. They further use permutation-based contrastive (hyvarinen2017nonlinear) to achieve identifiability on stationary time series data. The recent advancements in temporally causal representation include LEAP (yao2021learning), TDRL (yao2022temporally), which leverage the properties of independent noises and variability historical information. However, this assumption is too strong to meet in real-world scenarios since instantaneous dependencies are common in real-world scenarios like human motion. To solve this problem, researchers introduce further assumptions. For example, lippe2023causal propose the i-Citris, which demonstrates identifiability by assuming interventions on latent variables. In addition, morioka2023causal propose G-CaRL, which yields identifiability with the assumption of the grouping of observational variables. However, these assumptions are still hard to meet in practice. Hence we propose a more relaxed assumption, transition sparsity, to achieve identifiability under time series data with instantaneous
A.3 Causal Discovery with Latent Variables
@@
Several studies are proposed to discover causally related latent variables (chen2024identification; lachapelle2024additive; monti2020causal; zhang2012identifiability; tashiro2012estimation). Specifically, (huang2022latent; kong2024identification) leverage the vanishing Tetrad conditions or rank constraints to identify latent variables in linear-Gaussian models. xie2022identification; cai2019triad further draw upon non-Gaussianity in their analysis for linear, non-Gaussian scenarios. Other methods aim to reconstruct the hierarchical structures of the latent variables, like xie2022identification; huang2022latent. However, these methods usually use the linear assumption and can hardly handle the real-world time series data with complex nonlinear relationships.
A.4 Instantaneous Dependencey of Time Series Data
@@
When the discrete time series data is sampled in a low-frequency or low-resolution manner, instantaneous dependence (adak1996time; koutlis2019identification; gersch1985modeling; jamaludeen2022discovering; wu2023hierarchical) occurs where variables at each time step are not independent given their historical observations. swanson1997impulse first discuss the instantaneous dependence from the perspective of noise. Lots of works investigate instantaneous dependency from the causal view. Specifically, hyvarinen2010estimation estimate both instantaneous and lagged effects via a non-Gaussian model. gong2015discovering; gong2017causal discover causal structure with instantaneous dependency from subsampled and aggregated data. Recently, zhu2023beware consider instantaneous dependency in reinforcement learning.
A.5 Time Series Forecasting
@@
Time series forecasting (hyndman2018forecasting; box1970distribution) is one of the most popular research problems in recent years. Recently, the deep-learning-based methods have made great progress, which can be categorized according to different types of neural architectures like the RNN-based methods (hochreiter1997long; lai2018modeling; salinas2020deepar), CNN-based models (bai2018empirical; wang2022micn; wu2022timesnet), and the methods based on state-space model (gu2022parameterization; gu2021combining; gu2021efficiently). Recently, Transformer-based methods (zhou2021informer; wu2021autoformer; nie2022time) further push the development of time series forecasting. However, these methods seldom consider the instantaneous dependencies of time series data.
Appendix B Proof of Theory
B.1 Relationships between Ground-truth and Estimated Latent Variables
@@
(Identifiability of Temporally Latent Process) For a series of observations $\mathbf{x}_{t}\in\mathbb{R}^{n}$ and estimated latent variables $\mathbf{\hat{z}}_{t}\in\mathbb{R}^{n}$ with the corresponding process $\hat{f}_{i},\hat{p}(\epsilon),\hat{g}$, where $\hat{g}$ is invertible, suppose the process subject to observational equivalence $\mathbf{x}_{t}=\mathbf{\hat{g}}(\mathbf{\hat{z}}_{t})$. Let $\mathcal{M}_{\mathbf{z}_{t}}$ be the variable set of two consecutive timestamps and the corresponding Markov network respectively. Suppose the following assumptions hold:
  • A1 (Smooth and Positive Density): The conditional probability function of the latent variables ${\mathbf{c}}_{t}$ is smooth and positive, i.e., $p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})$ is third-order differentiable and $p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})>0$ over $\mathbb{R}^{2n}$,
@@
    $\begin{split}w(m)=&\Big(\frac{\partial^{3}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,1}^{2}\partial z_{t-1,m}},\cdots,\frac{\partial^{3}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,n}^{2}\partial z_{t-1,m}}\Big)\oplus\\ &\Big(\frac{\partial^{2}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,1}\partial z_{t-1,m}},\cdots,\frac{\partial^{2}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,n}\partial z_{t-1,m}}\Big)\oplus\Big(\frac{\partial^{3}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}\partial z_{t,j}\partial z_{t-1,m}}\Big)_{(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{z}}_{t}})},\end{split}$ (A1)
    where $\oplus$ denotes concatenation operation and $(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{z}}_{t}})$ denotes all pairwise indice such that $z_{t,i},z_{t,j}$ are adjacent in $\mathcal{M}_{{\mathbf{z}}_{t}}$. For $m\in[1,\cdots,n]$, there exist $4n+|\mathcal{M}_{{\mathbf{z}}_{t}}|$ different values of ${\mathbf{z}}_{t-1,m}$, such that the $4n+|\mathcal{M}_{{\mathbf{z}}_{t}}|$ values of vector functions $w(m)$ are linearly independent.
Then for any two different entries $\hat{z}_{t,k},\hat{z}_{t,l}$ of $\hat{{\mathbf{z}}}_{t}\in\mathbb{R}^{n}$ that are not adjacent in the Markov network $\mathcal{M}_{\hat{{\mathbf{z}}}_{t}}$ over estimated $\hat{{\mathbf{z}}}_{t}$,
(i) Each ground-truth latent variable $z_{t,i}$ of ${\mathbf{z}}_{t}\in\mathbb{R}^{n}$ is a function of at most one of $\hat{z}_{k}$ and $\hat{z}_{l}$,
(ii) For each pair of ground-truth latent variables $z_{t,i}$ and $z_{t,j}$ of ${\mathbf{z}}_{t}\in\mathbb{R}^{n}$ that are adjacent in $\mathcal{M}_{{\mathbf{z}}_{t}}$ over ${\mathbf{z}}_{t}$, they can not be a function of $\hat{z}_{t,k}$ and $\hat{z}_{t,l}$ respectively.
@@
$p(\hat{{\mathbf{x}}}_{t})=p({\mathbf{x}}_{t})\Longleftrightarrow p(\hat{g}(\hat{{\mathbf{z}}}_{t}))=p(g(z_{t}))\Longleftrightarrow p((g^{-1}\circ\hat{g})(\hat{{\mathbf{z}}}_{t}))=p({\mathbf{z}}_{t})\Longleftrightarrow p(h_{z}(\hat{{\mathbf{z}}}_{t}))=p({\mathbf{z}}_{t}),$ (A2)
where $\hat{g}:\mathcal{Z}\rightarrow\mathcal{X}$ denotes the estimated mixing function, and $h:=g^{-1}\circ\hat{g}$ is the transformation between the ground-truth latent variables and the estimated ones. Since $\hat{g}$ and $g$ are invertible, $h$ is invertible as well. Since Equation (A2) holds true for all time steps, there must exist an invertible function $h_{c}$ such that $p(h_{c}(\hat{{\mathbf{c}}}_{t}))=p({\mathbf{c}}_{t})$, whose Jacobian matrix at time step $t$ is
$\mathbf{J}_{h_{c},t}=\begin{bmatrix}\mathbf{J}_{h_{z},t-1}&0\\ 0&\mathbf{J}_{h_{z},t}\end{bmatrix}$. (A3)
@@
where * denotes any matrix. Since ${\mathbf{x}}_{t-1}$ can be fully characterized by itself, the left top and right top blocks are 1 and 0 respectively, and the determinant of this Jacobian matrix is the same as $|\mathbf{J}_{h_{z},t}|$. Therefore, we have:
$p(\hat{{\mathbf{z}}}_{t},{\mathbf{x}}_{t-1})=p({\mathbf{z}}_{t},{\mathbf{x}}_{t-1})|\mathbf{J}_{h_{z},t}|.$ (A4)
Dividing both sides of Equation (A4) by $p({\mathbf{x}}_{t-1})$, we further have:
$p(\hat{{\mathbf{z}}}_{t}|{\mathbf{x}}_{t-1})=p({\mathbf{z}}_{t}|{\mathbf{x}}_{t-1})|\mathbf{J}_{h_{z},t}|.$ (A5)
Since $p({\mathbf{z}}_{t}|{\mathbf{x}}_{t-1})=p({\mathbf{z}}_{t}|g({\mathbf{z}}_{t-1}))=p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})$, and similarly $p(\hat{{\mathbf{z}}}_{t}|{\mathbf{x}}_{t-1})=p(\hat{{\mathbf{z}}}_{t}|\hat{\mathbf{z}}_{t-1})$, we have:
$\log p(\hat{{\mathbf{z}}}_{t}|\hat{\mathbf{z}}_{t-1})=\log p({{\mathbf{z}}}_{t}|{\mathbf{z}}_{t-1})+\log|\mathbf{J}_{h_{z},t}|.$ (A6)